Afternoon - close loop with zero modes to verify communications. Slopes sent and wfs frame counter incrementing. Necesseary test after Adsec update made on 20210528 during the night for LBTI. OSCO to be performed later.

5:41 Chevron pattern on OCAM. Cycled but persisted. wfsinit. When WFS came up the adsec went down. Xianyu took a core dump. Restarted adsec.

Fully cycling the wfs again, still have the chevron

wfsinit the housekeeper did not come up cleanly, no wfs arbitrator so started those up through system processes.

6:13 Finally up and running. Sending OSCO script.

Estimated wavefront magnitude in AOS 8.70 - some light cloud wisps. DIMM 1.05". NCPA on.
TN#062024, 36, 48
TN#062306,18,30

-- JenniferPower - 29 May 2021
Topic revision: r1 - 29 May 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback